# A Comprehensive Evaluation of Large Language Models on Mental Illnesses

### v1

@@ 2.1 Experimental Setup
Our experiments were designed to assess LLM capabilities across three primary tasks: binary disorder detection, disorder severity evaluation, and psychiatric knowledge assessment. We conducted ZS experiments for all three tasks, where the LLM s were given only the input data (social media post or multiple-choice question) and the task description, without any prior examples. For the disorder severity prediction task, we also conducted FS experiments, where the LLM s were provided with three examples of posts and their corresponding severity ratings before being asked to predict the severity of a new post.
@@ 2.1.1 Task 1: ZS Binary Disorder Detection
2.1.2 Task 2: ZS/FS Disorder Severity evaluation
@@
2.1.3 Task 3: ZS Psychiatric Knowledge Assessment
@@ 2.2 Datasets
We considered various mental disorders but ultimately focused on depression, suicide risk, and stress due to their prevalence and the relative availability of high-quality datasets. Other disorders were not included in this study due to challenges in data collection and annotation. The sensitivity of mental health data often limits the public availability of relevant social media posts, and stricter data usage agreements imposed by social media platforms [^1] The landscape of social media data collection has changed significantly in recent years. In 2023, platforms like Reddit [22] and X (formerly Twitter) [23, 24] introduced substantial costs for accessing their API s, which were previously used to create many research datasets. Furthermore, stricter data usage agreements have been implemented, making some previously popular datasets, such as the popular SMHD [25], unsharable due to updated terms of service. have further restricted access to previously used datasets. Additionally, less prevalent disorders typically have fewer human-annotated datasets available, making it difficult to compile a sufficiently large and diverse dataset for robust model evaluation.
@@
Dreaddit [26] Binary Stress Classification Stress Reddit Crowdsourced 1000 posts Publicly Available
@@
DepSeverity [27] Depression Severity Classification Depression (4 levels) Reddit Professionals 3,553 posts Publicly Available
@@
Stress-Annotated- Dataset (SAD) [28] Stress Severity Classification Daily Stressors (9 categories) SMS-like conversations, crowdsourced data, LiveJournal Crowdsourced 6850 sentences Publicly Available
@@
DEPTWEET [29] Depression Severity Classification Depression (4 levels) X (formerly Twitter) Crowdworkers (trained) 40,191 tweets Available Upon Request
@@
CSSRS-Suicide [30] Suicide Risk Severity Assessment Suicide Risk (5 levels) Reddit (r/SuicideWatch) Professional psychiatrists 500 Reddit users Publicly Available
@@
SDCNL [31] Suicide vs. Depression Classification Depression, Suicidal Ideation Reddit (r/SuicideWatch, r/Depression) N/A 1,895 posts Publicly Available
@@
PsyQA [32] Generating Long Counseling Text Various Mental Health Disorders Mental Health Service Platform (Yixinli) Trained volunteers or professional counselors 22K questions, 56K answers Available Upon Request
@@
MedMCQA (Psychiatry Subset) [33] Medical Question Answering Various psychiatric disorders Medical entrance exams (AIIMS & NEET PG) Questions created by exam boards 4442 (~5%) Publicly Available
@@
PsySym [34] Symptom Identification 7 mental disorders (Depression, Anxiety, ADHD, Disorder, OCD , Eating Disorder) Reddit Trained volunteers and psychiatrists 8,554 sentences Available Upon Request
@@
MedQA (Psychiatry Subset) [35] Medical Question Answering Various clinical topics Medical entrance exams (AIIMS & NEET PG) Exam boards 47,457 Total, ~800 psychiatry related Publicly Available
@@
LGBTeen [36] Emotional Support and Information Queer-related topics Reddit (r/LGBTeens) Human Redditors and LLM s 1,000 posts, 11,320 LLM responses Publicly Available  [^2] While the author states that this dataset is available on Github, the dataset is not present in the Github repository provided by the author, and the author did not respond to our queries regarding the availability of the dataset
@@
ESConv [37] Emotional Support Conversation Various Mental Disorders Crowdsourced Crowdworkers (trained) 1,053 conversations Publicly Available
@@
Ji et al.[38] Suicidal Ideation and Mental Disorder Detection Suicidal Ideation, Depression, Anxiety, Bipolar Reddit (SuicideWatch and Mental Health-related subreddits) Experts 4800 posts Available Upon Request
@@
Cascalheira et al. [39] Gender Dysphoria Identification Gender Dysphoria Reddit (r/GenderDysphoria) Clinicians and Students with transgender and non-binary experience 1,573 posts Available Upon Request
@@
Hua et al. [40] Transgender and Gender Diverse Patient Identification Transgender and Gender Diverse Electronic Health Records (EHR) s Clinicians 3,350 patients Available upon request pending institutional reviews and approvals
@@
Mukherjee et al. [41] Autism Spectrum Disorder Symptom Identification Autism Spectrum Disorder Social networks (X), special children’s organizations Researchers Not Specified Not mentioned
@@
DATD [42] Depression and Anxiety Identification Depression and Anxiety X Not Specified 1,050 tweets Publicly Available
@@
PAN12 [43] Sexual Predator Identification Online Grooming Perverted Justice Website, Omegle, IRC logs One Expert Hundreds of thousands of conversations Publicly Available
@@
PJZC [44] Grooming Identification Online Grooming Perverted Justice Website, #ZIG IRC Channel, Chit-Chat Dataset Researchers 21,070 chats Publicly Available
@@
DAIC-WOZ [45] Distress Identification Depression, Post-traumatic stress disorder (PTSD), Anxiety Interviews Experts 421 interviews Available Upon Request
@@
D4 [46] Response Generation, Topic Prediction, Dialogue Summary, Severity Classification Depression, Suicide Risk Simulated conversations based on real patient portraits Psychiatrists, Crowdworkers 1,339 conversations Available Upon Request
@@
DialogueSafety [47] Dialogue Safety Classification Various (Nonsense, Humanoid Mimicry, Linguistic Neglect, etc.) Chinese Online Counseling Platform, Yixinli QA Column Experienced Supporters, Counseling Psychologists, Linguistic Experts 7,935 dialogues Available Upon Request
@@
CLPsych15 [48] Depression and PTSD Identification Depression, PTSD X Experts 1,746 tweets Available Upon Request pending institutional reviews and approvals
@@
RED SAM [49, 50] Depression Level Identification Depression (3 levels) Reddit Crowdworkers 16,632 posts Publicly Available
The datasets chosen for this study include:
2.2.1 Dreaddit [26]
@@
Dreaddit is a dataset for stress analysis collected from Reddit using the PRAW API. It contains 187,444 posts from 10 subreddits across five domains: interpersonal conflict (abuse, social), mental illness (anxiety, PTSD), and financial need (financial). For annotation, posts were divided into five-sentence segments, and crowdworkers on Amazon Mechanical Turk labeled them as "Stress," "Not Stress," or "Can’t Tell." Segments with low agreement or labeled as "Can’t Tell" were discarded, resulting in a final dataset of 3,553 labeled segments. The labels indicate whether the author of the segment is expressing stress and a negative attitude towards it.
2.2.2 DepSeverity [27]
@@ 2.2.3 SDCNL [31]
2.2.4 SAD [28]
SAD (Stress Annotated Dataset) is a dataset designed for the identification of daily stressors in SMS-like conversational systems. It contains 6,850 sentences labeled across nine categories of stressors, derived from stress management literature, live conversations from a prototype chatbot system, crowdsourced data, and targeted web scraping from LiveJournal. The dataset was created to improve chatbot interactions by enabling the identification and classification of stressors in text, facilitating appropriate interventions and empathetic responses. The stress categories include: Work (1043), School (1043), Financial Problem (152), Emotional Turmoil (238), Social Relationships (99), Family Issues (101), Health, Fatigue, or Physical Pain (103), Everyday Decision Making (101), and Other (123). However, in this analysis we focused on whether stress severity without considering the source of it.
@@ 2.2.4 SAD [28]
2.2.5 DEPTWEET [29]
DEPTWEET is a dataset designed for the detection of depression severity levels in social media texts. It contains 40,191 tweets labeled across four levels of depression severity: non-depressed, mild, moderate, and severe. The dataset was created by collecting tweets based on depression-related keywords and then having trained crowdworkers annotate them according to a typology based on the DSM-5 and Patient Health Questionnaire (PHQ)-9 clinical assessment tools.
@@ 2.2.5 DEPTWEET [29]
2.2.6 MedMCQA [33]
MedMCQA is a large-scale dataset designed for medical domain question answering, focusing on a wide range of medical subjects and topics relevant to medical entrance exams. It contains 194k high-quality multiple-choice questions (Multiple Choice Question (MCQ) s) covering 2.4k healthcare topics and 21 medical subjects, including a small portion (4442) related to psychiatry. Each sample in the dataset includes a question, correct answer(s), other options, and a detailed explanation of the solution. The dataset is designed to be challenging, requiring a deep understanding of both language and medical knowledge to answer the questions correctly. It is intended to serve as a benchmark for evaluating and improving the performance of AI models in medical question answering.
2.2.7 RED SAM [49, 50]
@@
2.2.8 CSSRS-Suicide (Excluded) [30]
CSSRS is a dataset designed for assessing the severity of suicide risk in individuals based on their social media posts. It contains 500 Reddit users from the r/SuicideWatch subreddit, labeled across five levels of suicide risk severity: Supportive, Indicator, Ideation, Behavior, and Attempt. The dataset was created to enable early intervention and prevention of suicide by identifying individuals who may be at risk.
@@ 2.2.8 CSSRS-Suicide (Excluded) [30]
We originally planned to include the CSSRS-Suicide dataset in our analysis, However, as the dataset is made up of posts from 500 people, with the number of posts included for each user varying greatly, from 1 in the lower case and more than 100 in the upper case. This made it difficult to evaluate the dataset fairly, with some users including more information than others. Furthermore, evaluating the users with too many posts would prove quite costly and exceed the context window of many of the smaller LLM s tested. Upon trying to include only one post per user (e.g. by selecting only the first post for each user), the results were significantly worse and as such the dataset was discarded.
@@ 2.3 Sampling
As all testing was conducted through API s, and for the bigger models, such as GPT-4, the costs could escalate rapidly, we employed two strategies for cost saving.First, we utilized the test dataset only if it was pre-split. Second, if the dataset exceeded 1000 examples, we employed a fair random sampling of 1000 instances. This sample size was chosen based on Hoeffding’s inequality (Equation 1) [51], which provides a general upper bound for the error difference between a sample and the entire population, given a random sample of that size.
@@
* $\nu$ is the empirical mean (sample average) * $\mu$ is the true mean (population average) * $\varepsilon$ is the deviation from the mean * $N$ is the number of samples
Plugging in the values $N$ = 1000 and $\varepsilon$ = 0.05 into Equation 1 (representing a maximum discrepancy of 5%), we calculate that the probability of the in-sample and out-of-sample error deviating by more than 5% is merely 0.0135. This implies a greater than 98.5% probability that the error will remain within a 5% range of the true error.

### v2

@@ 3.1 Experimental Setup
To ensure the robustness and interpretability of our results, we first conducted a performance variability experiment to evaluate the inherent performance fluctuations of each LLM. In this experiment, we assessed the inherent performance variability of each LLMon a representative task. The experiment involved repeatedly evaluating each model on the same dataset (DEPTWEET) five times using consistent, standard parameters (e.g., temperature 0). This approach allows us to approximate the natural performance fluctuations and establish a baseline for judging whether observed performance enhancements in the main tasks are meaningful improvements or simply due to random variability. The experiment was conducted on 1000 fairly sampled instances from the DEPTWEET dataset.
Following this initial variability assessment, our experiments were designed to assess LLMcapabilities across three primary tasks: binary disorder detection, disorder severity evaluation, and psychiatric knowledge assessment. We conducted ZSexperiments for all three tasks, where the LLMs were given only the input data (social media post or multiple-choice question) and the task description, without any prior examples. For the disorder severity prediction task, we also conducted FSexperiments, where the LLMs were provided with three examples of posts and their corresponding severity ratings before being asked to predict the severity of a new post.
@@ 3.1.1 Task 1: ZSBinary Disorder Detection
3.1.2 Task 2: FSDisorder Severity Evaluation
@@
3.1.3 Task 3: Psychiatric Knowledge Assessment
@@
3.1.4 Initial Exploration of LLMfine-tuning
While this study primarily focuses on evaluating the ZSand FScapabilities of LLMsin their pre-trained state, fine-tuning presents a complementary approach to further adapt these models for specific mental health tasks. To gain initial insights into this, we conducted preliminary fine-tuning experiments using GPT-4o mini and Phi-3.5-mini, employing parameter-efficient techniques like LoRA (Low-Rank Adaptation) for resource-conscious adaptation. For Phi-3.5-mini, we specifically utilized LoRA in 4-bit quantization (Q4) to further enhance efficiency.
In these preliminary fine-tuning experiments, we specifically fine-tuned both LLMsfor Task 2 (Disorder Severity Assessment) using the four severity-annotated datasets described in Section 3.2 (DepSeverity, DEPTWEET, RED SAM, and SAD). For GPT-4o-mini, fine-tuning was also initially explored on a 1600-sample balanced subset of the DEPTWEET dataset to assess in-domain performance. It is important to note that Phi-3.5-mini was released after the main zero-shot and few-shot experiments of this study were completed [^1] Consequently, Phi-3.5-mini is not included in the main zero-shot and few-shot analysis presented in Sections 4.. For all fine-tuning processes, we utilized fair random sampling on these datasets, adhering to the sampling strategy outlined in Section 3.3. We present and discuss the results of these initial fine-tuning experiments in Section 4.5.
@@ 3.2 Datasets
We considered various mental disorders but ultimately focused on depression, suicide risk, and stress due to their prevalence and the relative availability of high-quality datasets. Other disorders were not included in this study due to challenges in data collection and annotation. The sensitivity of mental health data often limits the public availability of relevant social media posts, and stricter data usage agreements imposed by social media platforms [^2] The landscape of social media data collection has changed significantly in recent years. In 2023, platforms like Reddit [Reddit_API_Changes_2023] and X (formerly Twitter) [Twitter_TOS_2023, Twitter_API_Changes_2023] introduced substantial costs for accessing their APIs, which were previously used to create many research datasets. Furthermore, stricter data usage agreements have been implemented, making some previously popular datasets, such as the popular SMHD [cohan2018smhd], unsharable due to updated terms of service. have further restricted access to previously used datasets. Additionally, less prevalent disorders typically have fewer human-annotated datasets available, making it difficult to compile a sufficiently large and diverse dataset for robust model evaluation.
@@
Table 1 summarizes the key characteristics of the datasets used in our experiments, including their clinical focus, task type, and scale. Below we provide detailed descriptions of these datasets, with particular attention to their annotation processes and label distributions.
Table 1: Summary of all the datasets used in this study.
Dataset Disorder Task Type Original Size Test Size
Dreaddit Stress Binary Classification 3,553 715
DepSeverity Depression Binary/Severity Classification 3,553 1,000
SDCNL Suicide Binary Classification 1,895 379
SAD Stress Binary/Severity Classification 6,850 1,000
DEPTWEET Depression Binary/Severity Classification 40,191 1,000
RED SAM Depression Binary/Severity Classification 16,632 1,000
MedMCQA Psychiatry Question Answering 4,442 1,000
Note: Test sizes reflect actual samples used after applying our sampling strategy (Section 3.3).
@@
Dreaddit [turcan2019dreaddit] Binary Stress Classification Stress Reddit Crowdsourced 1000 posts Publicly Available
@@
DepSeverity [naseem2022early] Depression Severity Classification Depression (4 levels) Reddit Professionals 3,553 posts Publicly Available
@@
Stress-Annotated- Dataset (SAD) [mauriello2021sad] Stress Severity Classification Daily Stressors (9 categories) SMS-like conversations, crowdsourced data, LiveJournal Crowdsourced 6850 sentences Publicly Available
@@
DEPTWEET [kabir2023deptweet] Depression Severity Classification Depression (4 levels) X (formerly Twitter) Crowdworkers (trained) 40,191 tweets Available Upon Request
@@
CSSRS-Suicide [gaur2019reddit] Suicide Risk Severity Assessment Suicide Risk (5 levels) Reddit (r/SuicideWatch) Professional psychiatrists 500 Reddit users Publicly Available
@@
SDCNL [haque2021deep] Suicide vs. Depression Classification Depression, Suicidal Ideation Reddit (r/SuicideWatch, r/Depression) N/A 1,895 posts Publicly Available
@@
PsyQA [sun2021psyqa] Generating Long Counseling Text Various Mental Health Disorders Mental Health Service Platform (Yixinli) Trained volunteers or professional counselors 22K questions, 56K answers Available Upon Request
@@
MedMCQA (Psychiatry Subset) [pal2022medmcqa] Medical Question Answering Various psychiatric disorders Medical entrance exams (AIIMS & NEET PG) Questions created by exam boards 4442 (~5%) Publicly Available
@@
PsySym [zhang2022symptom] Symptom Identification 7 mental disorders (Depression, Anxiety, ADHD, Disorder, OCD , Eating Disorder) Reddit Trained volunteers and psychiatrists 8,554 sentences Available Upon Request
@@
MedQA (Psychiatry Subset) [jin2021disease] Medical Question Answering Various clinical topics Medical entrance exams (AIIMS & NEET PG) Exam boards 47,457 Total, ~800 psychiatry related Publicly Available
@@
LGBTeen [lissak2024colorful] Emotional Support and Information Queer-related topics Reddit (r/LGBTeens) Human Redditors and LLMs 1,000 posts, 11,320 LLM responses Publicly Available  [^3] While the author states that this dataset is available on GitHub, the dataset is not present in the GitHub repository provided by the author, and the author did not respond to our queries regarding the availability of the dataset
@@
ESConv [liu2021towards] Emotional Support Conversation Various Mental Disorders Crowdsourced Crowdworkers (trained) 1,053 conversations Publicly Available
@@
Ji et al.[ji2022suicidal] Suicidal Ideation and Mental Disorder Detection Suicidal Ideation, Depression, Anxiety, Bipolar Reddit (SuicideWatch and Mental Health-related subreddits) Experts 4800 posts Available Upon Request
@@
Cascalheira et al. [cascalheira2023models] Gender Dysphoria Identification Gender Dysphoria Reddit (r/GenderDysphoria) Clinicians and Students with transgender and non-binary experience 1,573 posts Available Upon Request
@@
Hua et al. [hua2023deep] Transgender and Gender Diverse Patient Identification Transgender and Gender Diverse Electronic Health Records (EHR)s Clinicians 3,350 patients Available upon request pending institutional reviews and approvals
@@
Mukherjee et al. [mukherjee2023early] Autism Spectrum Disorder Symptom Identification Autism Spectrum Disorder Social networks (X), special children’s organizations Researchers Not Specified Not mentioned
@@
DATD [owen2020towards] Depression and Anxiety Identification Depression and Anxiety X Not Specified 1,050 tweets Publicly Available
@@
PAN12 [inches2012overview] Sexual Predator Identification Online Grooming Perverted Justice Website, Omegle, IRC logs One Expert Hundreds of thousands of conversations Publicly Available
@@
PJZC [milon2022take] Grooming Identification Online Grooming Perverted Justice Website, #ZIG IRC Channel, Chit-Chat Dataset Researchers 21,070 chats Publicly Available
@@
DAIC-WOZ [gratch2014distress] Distress Identification Depression, PTSD, Anxiety Interviews Experts 421 interviews Available Upon Request
@@
D4 [yao2022d4] Response Generation, Topic Prediction, Dialogue Summary, Severity Classification Depression, Suicide Risk Simulated conversations based on real patient portraits Psychiatrists, Crowdworkers 1,339 conversations Available Upon Request
@@
DialogueSafety [qiu2023benchmark] Dialogue Safety Classification Various (Nonsense, Humanoid Mimicry, Linguistic Neglect, etc.) Chinese Online Counseling Platform, Yixinli QA Column Experienced Supporters, Counseling Psychologists, Linguistic Experts 7,935 dialogues Available Upon Request
@@
CLPsych15 [coppersmith2015clpsych] Depression and PTSD Identification Depression, PTSD X Experts 1,746 tweets Available Upon Request pending institutional reviews and approvals
@@
RED SAM [sampath2022data, kayalvizhi2022findings] Depression Level Identification Depression (3 levels) Reddit Crowdworkers 16,632 posts Publicly Available
The following subsections detail the composition and annotation characteristics of each dataset:
3.2.1 Dreaddit [turcan2019dreaddit]
@@
Dreaddit is a dataset for stress analysis collected from Reddit using the PRAW API. It contains 187,444 posts from 10 subreddits across five domains: interpersonal conflict (abuse, social), mental illness (anxiety, Post-traumatic stress disorder (PTSD)), and financial need (financial). For annotation, posts were divided into five-sentence segments, and crowdworkers on Amazon Mechanical Turk labeled them as "Stress," "Not Stress," or "Can’t Tell." Segments with low agreement or labeled as "Can’t Tell" were discarded, resulting in a final dataset of 3,553 labeled segments. The labels indicate whether the author of the segment is expressing stress and a negative attitude towards it.
3.2.2 DepSeverity [naseem2022early]
@@ 3.2.3 SDCNL [haque2021deep]
3.2.4 SAD [mauriello2021sad]
SAD (Stress Annotated Dataset) is a dataset designed for the identification of daily stressors in SMS-like conversational systems. It contains 6,850 sentences labeled across nine categories of stressors, derived from stress management literature, live conversations from a prototype chatbot system, crowdsourced data, and targeted web scraping from LiveJournal. The dataset was created to improve chatbot interactions by enabling the identification and classification of stressors in text, facilitating appropriate interventions and empathetic responses. The stress categories include: Work (1043), School (1043), Financial Problem (152), Emotional Turmoil (238), Social Relationships (99), Family Issues (101), Health, Fatigue, or Physical Pain (103), Everyday Decision Making (101), and Other (123). However, in this analysis, we focused on the level of stress severity without considering its source.
@@ 3.2.4 SAD [mauriello2021sad]
3.2.5 DEPTWEET [kabir2023deptweet]
DEPTWEET is a dataset designed for the detection of depression severity levels in social media texts. It contains 40,191 tweets labeled across four levels of depression severity: non-depressed, mild, moderate, and severe. The dataset was created by collecting tweets based on depression-related keywords and then having trained crowdworkers annotate them according to a typology based on the DSM-Vand Patient Health Questionnaire (PHQ)-9 clinical assessment tools.
@@ 3.2.5 DEPTWEET [kabir2023deptweet]
3.2.6 MedMCQA [pal2022medmcqa]
MedMCQA is a large-scale dataset designed for medical domain question answering, and in this study, it was specifically used for the third task, psychiatric knowledge assessment (Section 4.4). focusing on a wide range of medical subjects and topics relevant to medical entrance exams. It contains 194k high-quality multiple-choice questions (Multiple Choice Question (MCQ)s) covering 2.4k healthcare topics and 21 medical subjects, including a small portion (4442) related to psychiatry. Each sample in the dataset includes a question, correct answer(s), other options, and a detailed explanation of the solution. The dataset is designed to be challenging, requiring a deep understanding of both language and medical knowledge to answer the questions correctly. It is intended to serve as a benchmark for evaluating and improving the performance of AI models in medical question answering.
3.2.7 RED SAM [sampath2022data, kayalvizhi2022findings]
@@
3.2.8 CSSRS-Suicide (Excluded) [gaur2019reddit]
CSSRS-Suicide is a dataset designed for assessing the severity of suicide risk in individuals based on their social media posts. It contains 500 Reddit users from the r/SuicideWatch subreddit, labeled across five levels of suicide risk severity: Supportive, Indicator, Ideation, Behavior, and Attempt. The dataset was created to enable early intervention and prevention of suicide by identifying individuals who may be at risk.
@@ 3.2.8 CSSRS-Suicide (Excluded) [gaur2019reddit]
We originally planned to include the CSSRS-Suicide dataset in our analysis, However, as the dataset is made up of posts from 500 people, with the number of posts included for each user varying greatly, from 1 in the lower end and more than 100 in the upper end. This made it difficult to evaluate the dataset fairly, with some users including more information than others. Furthermore, evaluating the users with too many posts would prove quite costly and exceed the context window of many of the smaller LLMs tested. Upon trying to include only one post per user (e.g. by selecting only the first post for each user), the results were significantly worse and as such the dataset was discarded.
@@ 3.3 Sampling
As all testing was conducted through APIs, and for the bigger models, such as GPT-4, the costs could escalate rapidly, we employed two strategies for cost saving. First, we utilized the test dataset only if it was pre-split. Second, if the dataset exceeded 1000 examples, we employed a fair random sampling of 1000 instances. This sample size was chosen based on Hoeffding’s inequality (Equation 1) [hoeffding1994probability], which provides a general upper bound for the probability that the difference between the empirical mean $\nu$ (i.e., the sample average) and the true mean $\mu$ (i.e., the population average) exceeds a predefined margin of error $\epsilon$, given a sample size NN
@@
Here, $N=1000$ represents the number of sampled instances, and $\epsilon=0.05$ corresponds to the maximum allowable deviation (5% discrepancy) between the sample and population metrics. Substituting these values into Equation 1, the probability of the empirical mean deviating from the true mean by more than 5% is bounded by $2e^{-2(0.05)^{2}\times 1000}\approx 0.0135.$ This implies a confidence level exceeding 98.65% that the observed error rate remains within 5% of the true error rate.